# Spatial Reasoning with Vision-Language Models in Ego-Centric Multi-View Scenes

### v1

@@ Abstract
Understanding 3D spatial relationships remains a major limitation of current Vision-Language Models (VLMs). Prior work has addressed this issue by creating spatial question-answering (QA) datasets based on single images or indoor videos. However, real-world embodied AI agents—such as robots and self-driving cars—typically rely on ego-centric, multi-view observations. To this end, we introduce Ego3D-Bench, a new benchmark designed to evaluate the spatial reasoning abilities of VLMs using ego-centric, multi-view outdoor data. Ego3D-Bench comprises over 8,600 QA pairs, created with significant involvement from human annotators to ensure quality and diversity. We benchmark 16 SOTA VLMs, including GPT-4o, Gemini1.5-Pro, InternVL3, and Qwen2.5-VL. Our results reveal a notable performance gap between human level scores and VLM performance, highlighting that current VLMs still fall short of human level spatial understanding. To bridge this gap, we propose Ego3D-VLM, a post-training framework that enhances 3D spatial reasoning of VLMs. Ego3D-VLM generates cognitive map based on estimated global 3D coordinates, resulting in 12% average improvement on multi-choice QA and 56% average improvement on absolute distance estimation. Ego3D-VLM is modular and can be integrated with any existing VLM. Together, Ego3D-Bench and Ego3D-VLM offer valuable tools for advancing toward human level spatial understanding in real-world, multi-view environments.
Figure 1: Ego3D-Bench is the first 3D spatial benchmark for VLMs using ego-centric multi-view images. It spans ego- and object-centric perspectives across 5 categories. A significant gap exists between human and VLM performance; our method, Ego3D-VLM, consistently narrows this gap.
@@ 1 Introduction
3D spatial understanding is a critical capability for embodied AI agents operating in the real world Chaplot et al. (2021); Xia et al. (2018); Wang et al. (2024b). This includes perceiving the location of surrounding objects, estimating their distances, and reasoning about their motion. VLMs have recently emerged as powerful tools to integrate visual perception and language reasoning, making them promising components for building intelligent embodied AI systems Ma et al. (2024b); Li et al. (2024b). Thus, enhancing and evaluating the spatial understanding abilities of VLMs has become an increasingly important research direction Team (2025); Cheng et al. (2024); Chen et al. (2024); Daxberger et al. (2025); Liao et al. (2024); Wang et al. (2024a); Wu et al. (2024b, a); Song et al. (2025); Ray et al. (2025).
@@
Recent benchmarks and datasets on spatial understanding primarily focus on spatial reasoning from single images Cheng et al. (2024); Chen et al. (2024); Liao et al. (2024) or videos captured in static, indoor environments Yang et al. (2024); Daxberger et al. (2025). In these cases, spatial understanding is framed around passive observations, where a single camera moves in a room to create a video, and the model is expected to infer spatial relationships in a relatively static scene. This setup differs fundamentally from the perceptual experience of real-world embodied agents, such as autonomous vehicles or mobile robots. These agents rely on ego-centric, multi-view observations Ma et al. (2024b), typically provided by multiple cameras simultaneously capturing the front, sides, and rear views of their surroundings. These views are not interchangeable or purely visual; they carry explicit spatial semantics tied to the agent’s frame of reference. For example, “left” and “right” refer to fixed directions relative to the agent’s body, and must be interpreted consistently over time as the agent moves through dynamic environments. This distinction is crucial: while prior video-based datasets may offer multiple viewpoints, they do not reflect the structured, directional, and temporally evolving nature of ego-centric multi-view inputs. Moreover, existing benchmarks do not evaluate the ability of VLMs to reason across these spatially grounded perspectives in dynamic, real-world scenes.
@@
This gap motivates the need for new benchmarks that better align with the spatial reasoning demands of embodied agents. To this end, we introduce Ego3D-Bench, a benchmark of 8,600 question-answer (QA) pairs carefully curated from the validation set of three public datasets: NuScenes Caesar et al. (2020), Waymo Open Dataset Sun et al. (2020), and Argoverse 1 Chang et al. (2019). Human annotators played a central role in both the dataset construction and the rigorous quality review process to ensure the reliability of the benchmark. We focused specifically on ego-centric multi-view tasks, rather than building a general-purpose benchmark, to complement existing monocular spatial benchmarks (e.g., VSI-Bench Yang et al. (2024)). As such, we excluded questions that can be answered by examining each view independently across multiple images (e.g., counting objects in each image) or by relying on general knowledge of LLMs (e.g., estimating the size of well-known objects). To our knowledge, Ego3D-Bench is the first benchmark to evaluate spatial reasoning of VLMs given ego-centric multi-view inputs.
@@
We evaluated 16 SOTA VLMs, including generalist and 3D spatial models on Ego3D-Bench, revealing a significant gap between human performance and current VLMs. We hypothesize that a key limitation lies in the inability of VLMs to construct a coherent world model from multi-view images. In contrast, humans naturally integrate visual information from their left, right, and front views into a unified spatial representation, enabling real-time reasoning and navigation. Prior work has attempted to bridge this gap by first generating a 3D point-cloud Team (2025); Huang et al. (2024); Hong et al. (2023b); Zhang et al. (2024a); Hong et al. (2023a) or rendering a bird-eye-view (BEV) image of the scene Qi et al. (2024). Although point-clouds and BEV images offer rich spatial information, they are challenging to reconstruct in dynamic environments, struggle with sparse multi-view inputs, and significantly increase inference time—often by a factor of ten Zhu et al. (2024).
@@
To this end, we propose, Ego3D-VLM, a post-training method that improves 3D spatial understanding of VLMs. The main idea of Ego3D-VLM is to create a textual cognitive map of the surrounding. The textual cognitive map defines a coordinate system center on the ego and locates important object (i.e., those referred to in the input prompt) in 3D coordinate space. Unlike point-clouds and BEV image methods Team (2025); Qi et al. (2024), our cognitive map only focuses on referred objects, making the number of input tokens significantly smaller and enabling efficient reasoning.
@@
Given multi-view images as input, we first use referring expression comprehension (REC) models to find the 2D location of referred expressions in pixel space. We also use a metric depth estimator to estimate the depth values. We then convert the 2D points to 3D points in camera coordinate space and transform 3D points from all views to the global coordinate space (i.e., front camera coordinate space). We define a cognitive map generator function that returns a textual cognitive map given 3D coordinates. The textual cognitive map defines a coordinate space for the VLM and organizes detected objects (expressions) based on the view-point. The textual cognitive map consistently improves SOTA VLMs on spatial reasoning.
In summary, our work has three main contributions:
@@
  • We introduce Ego3D-Bench, an ego-centric multi-view benchmark for evaluating the 3D spatial understanding of VLMs.
@@
  • We introduce Ego3D-VLM, a plug-and-play post-training framework to enhance 3D spatial understanding of VLMs, especially in ego-centric multi-view scenarios.
@@ 2 Related Work
3D Spatial Benchmarks for VLMs. VSI-Bench Yang et al. (2024), CA-VQA Daxberger et al. (2025), Q-Spatial-Bench Liao et al. (2024), SpatialRGPT-Bench Cheng et al. (2024), and All-Angle Bench Yeh et al. (2025) are five recent spatial benchmarks for VLMs. Q-Spatial-Bench and SpatialRGPT-Bench focus on single-view images. VSI-Bench and CA-VQA focus on video data captured from indoor static scenes. As noted, this setup differs fundamentally from the perceptual experience of real-world embodied agents, i.e., ego-centric multi-view images. All-Angle Bench is the first multi-view benchmark for VLMs. However, in their setup multiple cameras look at a scene from different direction—i.e., a setup for surveillance cameras and motion capture systems. Ego3D-Bench is the first ego-centric multi-view benchmark for VLMs created from dynamic outdoor scenes.
@@
The first group includes 3D-LLM Hong et al. (2023b), 3D-CLR Hong et al. (2023a), LEO Huang et al. (2024), SpatialLM Team (2025), and ChatScene Zhang et al. (2024a). While point-cloud representations offer rich spatial information, they are difficult to reconstruct in dynamic scenes, often struggle with sparse multi-view data, and significantly increase inference time (often by over 10×). The second group includes models such as LLaVA-3D Zhu et al. (2024), Video-3D LLM Zheng et al. (2024), GPT4Scene Qi et al. (2024), MM-Spatial Daxberger et al. (2025), SpatialVLM Chen et al. (2024), SpatialRGPT Cheng et al. (2024), and SpatialPIN Ma et al. (2024a). Our work falls in this image-based category.
LLaVA-3D and Video-3D LLM use depth maps and camera poses for 3D positional encoding. GPT4Scene leverages BEV to address 3D queries. However, these models are primarily trained on indoor, static scenes and are limited in handling quantitative spatial relationships such as object distances. SpatialVLM addresses this limitation by introducing a synthetic data generation pipeline, enabling more robust spatial reasoning remyxai (2024). SpatialRGPT enhances input representation by incorporating region proposals alongside the original image. MM-Spatial proposes a VLM that supports Chain-of-Thought spatial reasoning involving 2D grounding and depth estimation, and can also leverage depth input via tool-use. Different from SpatialRGPT, SpatialVLM, and MM-Spatial, our proposed Ego3D-VLM is a post-training method, can be applied to any existing VLM, and enhances spatial understanding of VLMs and outperforms prior works on ego-centric multi-view reasoning.
@@ 3 Ego3D-Bench
Ego3D-Bench is designed to quantitatively evaluate the 3D spatial understanding of VLMs from multi-view outdoor images. A major distinction between Ego3D-Bench and previous works is its focus on ego-centric multi-view data—an approach particularly relevant for applications in autonomous driving and robotics. Ego3D-Bench contains over 8,600 QA pairs in five distinct categories. This benchmark is constructed from the validation sets of three prominent outdoor multi-view datasets including nuScenes Caesar et al. (2020), Waymo Open Dataset Sun et al. (2020), and Argoverse 1 Chang et al. (2019), featuring 6, 7, and 5 distinct camera views, respectively (Figure 2). These datasets cover diverse outdoor environments, including urban areas, highways, and rural regions. Ego3D-Bench leverages the multi-view nature of these datasets to formulate questions that require fine-grained visual comprehension across different viewpoints as well as reasoning over 3D spatial relationships. In this section, we describe the detailed process of constructing the questions of Ego3D-Bench (Figure 2).
Figure 2: Overview of our benchmark creation pipeline. Human annotators played a key role throughout the process. On the right, we show the distribution of the samples in our Ego3D-Bench. Ego.: Ego-centric, Obj.: Object-centric.
@@ 3.1 Benchmark Construction
Creation of Source Files: Outdoor datasets such as the ones used in this work, unlike their indoor counterparts, often feature multiple instances of the same object within each scene. While indoor environments typically contain unique items—such as a single oven or television per room—outdoor scenes commonly include numerous similar objects, like multiple cars or pedestrians. This makes it challenging to uniquely reference a target object in the scene. Thus, annotators begin by carefully reviewing each scene to identify unique objects (called "Filtering" in Figure 2). They then compose concise captions describing these objects (called "Captioning" in Figure 2). The descriptions are designed to be short, yet discriminative such that each object can be uniquely identified. Furthermore, the ground-truth 3D annotations (bounding boxes) are collected from the source datasets. Object captions, 3D annotations, and the camera view from which the object is visible are used to construct a source file for QA creation.
Creation of QAs: Each question category follows a predefined template with placeholders for object names, such as How far is <obj1> in <view1> from <obj2> in <view2>? (see Appendix for all question templates). Each question is constructed by placing the generated object annotation and camera views from the source file in the question template. To generate the answers, we use rule-based functions that leverage ground-truth 3D annotations. For challenging categories, such as motion reasoning, a final human review is conducted to ensure the accuracy of the QA pairs.
@@ 3.2 Benchmark Details
Figure 3 illustrates different question categories in Ego3D-Bench. We emphasize on questions that require understanding the 3D world space by analyzing and consolidating information from multiple views. Thus, we exclude questions that can be answered using a single-view or by analyzing each view independently (e.g., counting the number of objects), as they do not contribute to multi-view spatial reasoning. We define question from the ego-perspective and from the perspective of objects in the scene. To clearly indicate the perspective of each question, we categorize them into ego-centric or object-centric. In the following, we describe the five tasks (each composed of ego-centric and object-centric types) in our benchmark.
(1) Absolute Distance Measurement. This category asks the VLM to estimate the absolute metric distance between the ego and another object in the scene or between two different objects from different camera views. As shown in Figure 3, this category is designed in two forms of multi-choice QA and absolute meter.
@@ 4 Post-Training 3D Spatial Understanding: Ego3D-VLM
Camera Coordinate Transformation. To obtain 3D spatial information, we use a metric depth estimator to predict a dense depth map $D^{(v)}\in\mathbb{R}^{H\times W}$ for each view $v$. For each detected object center $\mathbf{u}_{i}^{(v)}=(x_{i},y_{i})^{\top}$, we extract the corresponding depth value $d_{i}^{(v)}=D^{(v)}(x_{i},y_{i})$. We then project each center point into the 3D camera coordinate system using the camera intrinsics $K^{(v)}\in\mathbb{R}^{3\times 3}$:
@@
Creating a Cognitive Map. We define a cognitive map generator function, $\mathcal{F}_{\text{cog}}$, which takes as input the set of 3D global coordinates and corresponding referring expressions for all detected objects across all camera views. Specifically, for each object $i$ detected from view $v$, we denote its global 3D position as $\mathbf{p}_{\text{global},i}^{(v)}$ and its matched referring expression as $c_{i}^{(v)}$. The function outputs a textual cognitive map $C$, defined as:
@@
Given a VLM as $\mathcal{V}$ that answers queries using visual and textual contexts, it takes as input the cognitive map $C$, a set of multi-view images $\mathcal{I}=\{I^{(v)}\}_{v}$, and a natural language query $q$, and returns an answer $a$:

### v2

@@ Abstract
Understanding 3D spatial relationships remains a major limitation of current Vision-Language Models (VLMs). Prior work has addressed this issue by creating spatial question-answering (QA) datasets based on single images or indoor videos. However, real-world embodied AI agents—such as robots and self-driving cars—typically rely on ego-centric, multi-view observations. To this end, we introduce Ego3D-Bench, a new benchmark designed to evaluate the spatial reasoning abilities of VLMs using ego-centric, multi-view outdoor data. Ego3D-Bench comprises over 8,600 QA pairs, created with significant involvement from human annotators to ensure quality and diversity. We benchmark 16 SOTA VLMs, including GPT-4o, Gemini1.5-Pro, InternVL3, and Qwen2.5-VL. Our results reveal a notable performance gap between human level scores and VLM performance, highlighting that current VLMs still fall short of human level spatial understanding. To bridge this gap, we propose Ego3D-VLM, a post-training framework that enhances 3D spatial reasoning of VLMs. Ego3D-VLM generates cognitive map based on estimated global 3D coordinates, resulting in 12% and 56% average improvements on multi-choice QA and absolute distance estimation, respectively. Ego3D-VLM can be integrated with any existing VLM. Together, Ego3D-Bench and Ego3D-VLM offer valuable tools for advancing toward human level spatial understanding in real-world, multi-view environments.
Figure 1: Ego3D-Bench is a 3D spatial benchmark for VLMs using ego-centric multi-view images. It spans ego- and object-centric perspectives across 5 categories. A significant gap exists between human and VLM performance; our method, Ego3D-VLM, consistently narrows this gap.
@@ 1 Introduction
3D spatial understanding is a critical capability for embodied AI agents operating in the real world [4, 40]. This includes perceiving the location of surrounding objects, estimating their distances, and reasoning about their motion. VLMs have recently emerged as powerful tools to integrate visual perception and language reasoning, making them promising components for building intelligent embodied AI systems [28, 23]. Thus, enhancing and evaluating the spatial understanding of VLMs has become an increasingly important research direction [38, 6, 5, 51, 50, 39, 42, 41, 36, 33].
@@
Recent benchmarks on spatial understanding focus mainly on spatial reasoning from single images [6, 5, 24] or videos captured in static indoor environments [44, 8]. In such cases, spatial understanding is framed around passive observations, where a single camera moves in a room to create a video, and the model should infer spatial relationships in a relatively static scene. This setup differs fundamentally from the perceptual experience of real-world embodied agents such as self-driving cars or mobile robots. These agents rely on ego-centric multi-view observations [28], provided by multiple cameras simultaneously capturing front, side, and rear views of their surroundings. These views are not interchangeable or purely visual; they carry explicit spatial semantics tied to the agent’s frame of reference. E.g, “left” & “right” refer to fixed directions relative to the agent’s body, and must be interpreted consistently over time as the agent moves through dynamic environments. This distinction is crucial: while prior video-based datasets may offer multiple viewpoints, they do not reflect the structured, directional, and temporally evolving nature of ego-centric multi-view inputs. Also, existing benchmarks do not evaluate VLMs’ reasoning ability across these spatially grounded perspectives in dynamic, real-world scenes.
@@
This gap motivates the need for new benchmarks that better align with the spatial reasoning demands of embodied agents. To this end, we introduce Ego3D-Bench, a benchmark of 8.6K QA pairs carefully curated from the validation set of 3 public datasets: NuScenes [2], Waymo Open Dataset [37], and Argoverse 1 [3]. Human annotators played a central role in both the dataset construction and the rigorous quality review process to ensure the reliability of the benchmark. We focused specifically on ego-centric multi-view tasks, rather than building a general-purpose benchmark, to complement existing monocular spatial benchmarks (e.g., VSI-Bench [44]). Thus, we excluded questions that can be answered based on each view independently across multiple images (e.g., counting objects in each image) or general knowledge of LLMs (e.g., estimating the size of well-known objects). To our knowledge, Ego3D-Bench is the first benchmark to evaluate spatial reasoning of VLMs given ego-centric multi-view inputs.
@@
We evaluated 16 SOTA VLMs, including generalist and 3D spatial ones on Ego3D-Bench, revealing a significant gap between human performance and current VLMs. We hypothesize that a key limitation lies in the inability of VLMs to construct a coherent world model from multi-view images. In contrast, humans naturally integrate visual info from their left, right, and front views into a unified spatial representation, enabling real-time reasoning and navigation. Prior work has attempted to bridge this gap by first generating a 3D point-cloud [38, 19, 18, 47, 17] or rendering a bird-eye-view (BEV) image of the scene [31]. Although point-clouds and BEV images offer rich spatial information, they are challenging to reconstruct in dynamic environments, struggle with sparse multi-view inputs, and significantly increase inference time—often by a factor of ten [53].
@@
To this end, we propose, Ego3D-VLM, a post-training method that improves 3D spatial understanding of VLMs. The main idea of Ego3D-VLM is to create a textual cognitive map of the surrounding. The textual cognitive map defines a coordinate system center on the ego and locates important object (i.e., those referred to in the input prompt) in 3D coordinate space. Unlike point-clouds and BEV image methods [38, 31], our cognitive map only focuses on referred objects, making the number of input tokens significantly smaller and enabling efficient reasoning.
@@
Given multi-view images as input, we first use referring expression comprehension (REC) models to find the 2D location of referred expressions in pixel space. We also use a metric depth estimator to estimate the depth values. We then convert the 2D points to 3D points in camera coordinate space and transform 3D points from all views to the global coordinate space (i.e., front camera coordinate space). We define a cognitive map generator function that returns a textual cognitive map given 3D coordinates. The textual cognitive map defines a coordinate space for the VLM and organizes detected objects (expressions) based on the view-point. The textual cognitive map consistently improves SOTA VLMs on spatial reasoning. The major contributions of our work are:
@@
  • We introduce Ego3D-Bench, an ego-centric multi-view benchmark to evaluate 3D spatial understanding of VLMs.
@@
  • We propose Ego3D-VLM, a plug-and-play post-training framework to enhance 3D spatial understanding of VLMs, especially in ego-centric multi-view scenarios.
@@ 2 Related Work
3D Spatial Benchmarks/Datasets for VLMs. We categorize prior datasets and benchmarks into three groups based on the input data type. (1) [25, 20, 24, 36, 50, 6, 10] focus on single-view images. (2) [44, 8, 22, 45, 33] focus on videos captured from indoor static scenes. As noted, this setup differs from the perceptual experience of real-world embodied agents, which involves ego-centric multi-view images. (3) All-Angle [46] is the first multi-view benchmark for VLMs. However, in their setup, multiple cameras observe a scene from different directions—similar to surveillance or motion capture systems. In contrast, Ego3D-Bench is the first ego-centric multi-view benchmark for VLMs created from dynamic outdoor scenes.
Figure 2: Overview of our Ego3D-Bench creation pipeline. Human annotators played a key role in the process. Right: the distribution of the samples. Ego.: Ego-centric, Obj.: Object-centric.
@@
[18, 17, 19, 38, 49, 47] fall in the first group. While point-cloud representations offer rich spatial information, they are difficult to reconstruct in dynamic scenes, often struggle with sparse multi-view data, and significantly increase inference time by over 10×. The 2nd group includes LLaVA-3D [53], Video-3D LLM [52], GPT4Scene [31], MM-Spatial [8], SpatialVLM [5], SpatialRGPT [6], and SpatialPIN [27]. Our work falls in this image-based category.
LLaVA-3D and Video-3D LLM use depth maps and camera poses for 3D positional encoding. GPT4Scene leverages BEV to address 3D queries. However, these models are primarily trained on indoor, static scenes and limited in handling quantitative spatial relationships such as object distances. SpatialVLM addresses this limitation by introducing a synthetic data generation pipeline, enabling more robust spatial reasoning [34]. SpatialRGPT enhances input representation using region proposals alongside the original image. MM-Spatial proposes a VLM that supports Chain-of-Thought spatial reasoning involving 2D grounding and depth estimation, and can also leverage depth input via tool-use. Different from SpatialRGPT, SpatialVLM, and MM-Spatial, our proposed Ego3D-VLM is a post-training method, can be applied to any existing VLM, and enhances spatial understanding of VLMs and outperforms prior works on ego-centric multi-view reasoning.
@@ 3 Ego3D-Bench
Ego3D-Bench is designed to quantitatively evaluate 3D spatial understanding of VLMs from multi-view outdoor images. A major distinction between Ego3D-Bench and previous works is its focus on ego-centric multi-view data—an approach particularly relevant for applications in self-driving and robotics. Ego3D-Bench contains over 8.6K QA pairs in 5 distinct categories. This benchmark is constructed from the validation sets of 3 prominent outdoor multi-view datasets: nuScenes [2], Waymo Open Dataset [37], and Argoverse 1 [3], featuring 6, 7, and 5 distinct camera views, respectively (Figure 2). These datasets cover diverse outdoor environments, including urban areas, highways, and rural regions. Ego3D-Bench leverages the multi-view nature of these datasets to formulate questions that require fine-grained visual comprehension across different viewpoints as well as reasoning over 3D spatial relationships. In this section, we describe the detailed process of constructing the questions of Ego3D-Bench.
@@ 3.1 Benchmark Construction
Creation of Source Files: Outdoor datasets, unlike their indoor counterparts, often feature multiple instances of the same object within each scene. While indoor environments typically contain unique items—such as a single oven or television per room—outdoor scenes commonly include numerous similar objects, like multiple cars or pedestrians. This makes it challenging to uniquely reference a target object in the scene. Thus, annotators begin by carefully reviewing each scene to identify unique objects (called "Filtering" in Figure 2). They then compose concise captions describing these objects (called "Captioning" in Figure 2). The descriptions are designed to be short, yet discriminative such that each object can be uniquely identified. Furthermore, the ground-truth 3D annotations (bounding boxes) are collected from the source datasets. Object captions, 3D annotations, and the camera view from which the object is visible are used to construct a source file for QA creation.
Creation of QAs: Each question category follows a predefined template with placeholders for object names, such as How far is <obj1>in <view1>from <obj2>in <view2>? (see Appendix A.2 for all question templates). Each question is constructed by placing the generated object annotation and camera views from the source file in the question template. To generate the answers, we use rule-based functions that leverage ground-truth 3D annotations. For challenging categories, such as motion reasoning, a final human review is conducted to ensure the accuracy of the QA pairs.
@@ 3.2 Benchmark Details
Figure 3 illustrates different question categories in Ego3D-Bench. We emphasize on questions that require understanding the 3D world by utilizing information from multiple views. Thus, we exclude questions that can be answered using a single-view or by analyzing each view independently (e.g., counting the number of objects), as they do not contribute to multi-view spatial reasoning. We define question from the perspective of both ego and objects in the scene. To clearly indicate the perspective of each question, we categorize them into ego-centric or object-centric. In the following, we describe the five tasks (each composed of ego-centric and object-centric types) in our benchmark.
(1) Absolute Distance Measurement. This category asks the VLM to estimate the absolute metric distance between the ego and another object in the scene or between two different objects from different views. This category is designed in two forms of multi-choice QA and absolute meter.
@@ 4 Post-Training 3D spatial understanding : Ego3D-VLM
Camera Coordinate Transformation. To obtain 3D spatial info, we use a metric depth estimator to predict a dense depth map $D^{(v)}\in\mathbb{R}^{H\times W}$ for each view $v$. For each detected object center $\mathbf{u}_{i}^{(v)}=(x_{i},y_{i})^{\top}$, we extract the corresponding depth value $d_{i}^{(v)}=D^{(v)}(x_{i},y_{i})$. We then project each center point into the 3D camera coordinate system using the camera intrinsics $K^{(v)}\in\mathbb{R}^{3\times 3}$:
@@
Creating a Cognitive Map. We define a cognitive map generator function $\mathcal{F}_{\text{cog}}$, which takes as input the set of 3D global coordinates and corresponding referring expressions for all detected objects across all camera views. Specifically, for each object $i$ detected from view $v$, we denote its global 3D position as $\mathbf{p}_{\text{global},i}^{(v)}$ and its matched referring expression as $c_{i}^{(v)}$. The function outputs a textual cognitive map $C$, defined as:
@@
Given a VLM $\mathcal{V}$ that answers queries by visual-textual contexts, it takes as input the cognitive map $C$, a set of multi-view images $\mathcal{I}=\{I^{(v)}\}_{v}$, and a natural language query $q$, to return an answer $a$: